1. Below is a plot of the survival function for a random variable X 2 (a) What properties of this graph guarantee that it is a survival function for a discrete random variable? (b) Find P[x -3) (c) Find P[X > 3) (d) Find P[X <3 (c) lind the probability, maz彰fractiou for K
Answer #1

a)

for the distribution is decreasng which indicate it is a survivial distribution funtion

also as CDF goes from 0 to 1 which fullfills valid probability distribution condition

and being step wise function denotes it is discrete,

b)

P(X=3)=0.8-0.45=0.35

c)P(X>3)=0.45

d) P((X<3)=0.80

e)below is pmf of X:

P(X=1)=1-0.85 =0.15

P(X=2)=0.85-0.8=0.05

P(X=3)=0.8-0.45=0.35

P(X=5)=0.45-0.25=0.20

P(X+6)=0.25-0 =0.25